•Merix Corp. is considering issuing $500 million of additional debt and using the proceed to buy back some outstanding shares.

•Merix shares are currently trading at $45 per share and there are 30 million shares outstanding.

•Merix also has $850 million of debt outstanding which is trading at par at a yield of 9.5%.

•Analysts are forecasting that Merix shares will pay a dividend of $1.20 in one year and that the dividend will grow at 10% per year for the foreseeable future.

•Merix has a corp. tax rate of 46%

a) What is the WACC assuming Merix’s current capital structure?

b) Find the value of the the equivalent unlevered firm and the required return on unlevered’s assets.

c) Find the value of Merix if it undertakes the restructuring. Find the new value of the equity.

d) What is the new value of the shares and how many shares could be repurchased?

e) Find the WACC of Merix under the new capital structure.
Answer #1


V= D + E + P(ignore)

E (common shares): 30000000 shares x 45= 1350000000

D(bounds): 850000000, this is trading as par/ book/face value, market rate = book value

V= 1350000000 + 850000000= 2200000000

#2Find the value of the equivalent unlevered firm and the required return on unlevered assets, find the cost return

Rd =pretax cost of debt (YTM)=9.5%,Make after tax, After Tax= Pretax x (1-t)= 0.095 x (1-0.46)= 0.0513

Re= Cost of Equity = (D1/Po)+g where D1=nest years dividend, Po= Current share price, g=growth rate

Re= (1.20/45) + 0.10= 0.1267

List1 (value weights): 1350000000 (E), 850000000 (D)

List 2 (return): 0.1267 (Re), 0.0513 (Rd)(1-t)Calc, 2-var, X= WACC= 0.097568, which is 9.75%

To find weight of equity, E/V= 1350000000/ 220000000

Thanks